OpenStudy (nannie):

HELP! A rectangle's area is A=4²ft. and the width is W=L-3ft. What is the length(L) and width (W). Please show your work so I will understand how you got the answer. Thank you!

OpenStudy (nannie):

The area is 4^2

OpenStudy (luffingsails):

gah. sorry missed that. I'll delete and repost.

OpenStudy (luffingsails):

Method is the same though, just change the numbers.

OpenStudy (nannie):

I had (l=7)(L=-1) Didn't get the method

OpenStudy (luffingsails):

Are you sure it doesn't say: \[4ft^2\] ? Because that is not the same as \[4^2\]

OpenStudy (nannie):

yes it did!

OpenStudy (luffingsails):

Okay, so what I posted originally is right: Given A = 4ft^2 A = L * W W = L - 3 So: \[L^2-3L-4=0\] \[(L-4)(L+1)=0\] So, L = 4 or -1 Length can't be negative, so L = 4. W = L - 3 = 4 - 3 = 1. Make sense?

OpenStudy (nannie):

Yes, Thanks!

OpenStudy (luffingsails):

Cool.